OpenStudy (anonymous):

WIN MEDAL!! A cylinder and a rectangular prism have the same volume and same height. The base of the prism is a square with side length 8 cm. What is the approximate radius of the cylinder?

OpenStudy (anonymous):

WIN MEDAL!!

OpenStudy (anonymous):

Pleeaassee hellppp!!

OpenStudy (anonymous):

Let the height of the cylinder and the prism be h, radius of the cylinder be r. Volume of the cylinder = pi * r^ 2 *h Volume of the prism = 8^2 * h Since both the volumes are equal pi*r^2*h = 8^2 *h r=8/root pi r = 4.52 cm

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

:)